OpenStudy (anonymous):

solve on interval 0,2pi

OpenStudy (anonymous):

\[1+\sin \theta = \frac{ \sqrt{3}+2 }{ 2 }\]

OpenStudy (jdoe0001):

so, if you solve for \(sin(\theta) \) what does it give you?

OpenStudy (anonymous):